OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4

OpenStudy (anonymous):

3(x + 1) = -2(x - 1) - 4 - distribute the 3 and -2 3x+3= -2x+2-4 --- combine like terms 3x+3 = -2x-2-- subtract 3 to both sides and add 2x to both sides 5x = -5 -- divide both sides by 5 x = -1